OpenStudy (anonymous):

36% of what number is 18? How do I solve this?

OpenStudy (anonymous):

let the number be X 36X/100 = 18 X = (18 * 100) / 36 X = 50.

OpenStudy (anonymous):

thanks

OpenStudy (skullpatrol):